# access point ADAS advanced driver-assistance systems

### v1

@@ II-D Radar MAC Protocol
In dense vehicular networks, in particular where the radar and communication bands are shared, frequent radar signals may degrade the overall detection as well as the communication performance. For this, we propose a tunable block ALOHA protocol for shared radar channel access. Let the time frames be divided into blocks of $T$ slots each. In each block, only a subset of the vehicles are allowed to activate their radars. The activation set follows the classical ALOHA protocol with parameter $\delta\in[0,1]$. Thus, each vehicle in $\Phi_{\rm P}$ operates in a block with probability $\delta$. Within the block, it attempts to detect the target in each of the $T$ slots. We present our analysis from the perspective of a single such block in which the ego radar is active. Later, we discuss the intricacies of designing the block length in Section IV. Here we define the key metrics that we statistically characterize in the next section.
@@ III Statistics of Radar Tracking
The conditional success probability in (2) is different from that derived in [5, 10, 11]. In particular, the above works consider random access in each slot and hence, there the transmission state of each radar needs to be averaged out with respect to the ALOHA parameter. However, in our case, once the transmission states of the radars are determined at the beginning of the block, the same remains fixed throughout all the slots of the block. Hence, the impact of the ALOHA only appears in the set over which the product of the point functions is evaluated in (2).
@@
Remark 2.
Leveraging de Moivre’s solution [23, Section 22.6], the conditional complementary cumulative density function (CCDF) ${F}_{L}(\nu)=\mathbb{P}\left(L\geq\nu\mid\Phi\right)$ of $L$ as a function of $P_{\rm d}$ is
@@
Given that the ego radar is active in a block, the moments of conditional success probability $P_{\rm d}$ of the typical controller at a given time $t$ within the block is given by
@@
$=\exp\left(-\lambda\delta\int_{0}^{\infty}1-\mathcal{T}_{z}^{l}{\rm d}z\right)\cdot\exp\left(-2\lambda_{L}\int_{0}^{2\pi}\int_{0}^{\infty}1-\right.$
$\left.\exp\left(-\delta\lambda\int_{a_{i}}^{b_{i}}1-\mathcal{T}^{l}_{z,x}{\rm d}z\right){\rm d}x{\rm d}\theta\right),$ (4)
where $\mathcal{T}_{z}=\left(\frac{R^{-2\alpha}}{R^{-2\alpha}+\gamma z^{-\alpha}}\right)$, $\mathcal{T}_{z,x}=\frac{R^{-2\alpha}}{R^{-2\alpha}+\gamma w(x,z,\theta))^{-\alpha}}$ and $w(x,z,\theta)=x^{2}+z^{2}+2xz\cos\theta$.
@@ IV QoS-Aware Optimal Frame Design
In this section, we use the above framework for the design of the transmit frames and the selection of optimal access probability from the perspective of specific automotive use-cases. Let us define $T_{\rm L}$ as the latency requirement specified by the use-case. As an example, low-latency applications such as lane change assistance and automatic emergency braking (AEB) may employ a $T_{\rm L}$ of the order of 1 ms (note that this not the end-to-end latency, but only the latency over the radio link). On the contrary, use-cases that have less stringent latency requirements such as blind spot detection, park assist etc have $T_{\rm L}$ of the order of 10 ms. Some indicative latency requirements are outlined in Table I. Furthermore, we introduce a reliability measure based on the parameter $\nu$. in particular, a larger value of $\nu$ corresponds to a higher reliability requirement from the service. For the tracking service to be rendered successful, let us assume that within the latency deadline, the use-case necessitates at least one successful tracking event, i.e., the ego radar experiences at least one burst of $\nu$ consecutive successes within a deadline of $T_{\rm L}$.
@@ V Numerical Results and Discussion
In this section we discuss the salient features of the network with the help of some numerical results. The radar parameters used for the numerical results are taken from [24]. Specifically, $P=10$ dBm, $\bar{\sigma}=30$ dBsm, $\alpha=2$, $G_{\rm t}=10$ dBi, $f_{c}=76.5$ GHz, and $N_{0}=-174$ dBm/Hz. Unless otherwise stated, we assume $R=15$ m. The street and vehicle densities are from [17] where the authors have reported real-world data on the same.
@@ V-A Detection vs tracking performance
V-B Impact of Street and Vehicle density
@@
Figure 4: Expected number of bursts with a length $\geq\nu=8$ for $T=20$ for different $\lambda_{P}$ and $\lambda_{L}$. Here we have assumed
@@
TABLE I: Use cases and corresponding latency and reliability requirement derived from [25] and [26]. However, note that the latency requirement mentioned here is not end-to-end latency, but the latency over the radio link. The figures are indicative and real values are system dependent. The optimal channel access probability is calculated against each use-case. We assume $\lambda=5e-2$ m-1 and a beamwidth of 10 degrees.
V-C Impact of Use-Case QoS Requirements
@@
Fig. 5 shows the impact of the latency deadline $T_{\rm L}$ and the reliability requirement modeled as the number of consecutive slots of detection $\nu$. Specifically, we study two services - lane change assist [27] with a typical latency requirement of the order of 10 ms, and automatic emergency braking [28] with a typical latency of 1-3 ms. Note that these latency requirements from the radio link and not end-to-end latency. Furthermore, let us specify two reliability requirements: $\nu=8$ and $\nu=12$. For the results of this figure, we fix the number of slots per block as $T=20$. We assume a pulse width of $200$ $\mu$s. This results in $K=50$ blocks for the lane change assist use-case and $k=15$ blocks for the automatic emergency braking use-case. Recall that for the service to be rendered successful, the ego radar needs at least one successful tracking event. Due the larger number of blocks to compete in for access, the success probability is higher for the lane change assist use-case as compared to the automatic emergency braking use-case. Interestingly, for higher reliability requirement ($\nu=12$), not only the success probability is lower, but also the optimal $\delta$ is higher. In Table I we show example use-cases with corresponding latency and reliability requirements. Accordingly, the optimal channel access probability is demonstrated. The optimization of $\delta$ is not straightforward due to the two competing phenomenon of interference and channel access. in this regard, we discuss the considerations for the optimal frame design next.
V-D Optimal Frame Design
Figure 6: Impact of the frame design policy on the tracking performance. Here we use real data fitting of New Delhi city to extract the PLCP parameter $\lambda_{L}=0.004$ m-1 (see [17] for details) and use a vehicle density of $\lambda=5e-2$ m-1 and $\Omega=10$ degrees.
@@ Appendix A Proof of Lemma 1
We derive the expressions of $a$ and $b$ using simple geometrical arguments from Fig. 8 that shows the ego radar in green and an interfering radar in red. Here the point O is the origin and the y-axis represents the street containing the ego radar. The affine extension of the segment BE represents an intersecting line containing the interfering radar. This line intersects $y$-axis at a distance $d$ from the origin. The interfering radar is present at the location $a$, i.e., the minimum distance along GB for a radar to cause interference. The corresponding maximum distance is represented as $b$.

### v2

@@ II-D Radar MAC Protocol
Figure 2: Illustration of the frame design and the MAC protocol from the perspective of the ego-radar.
In dense vehicular networks, in particular where the radar and communication bands are shared, frequent radar signals may degrade the overall detection as well as the communication performance. For this, we propose a tunable block ALOHA protocol for shared radar channel access. We adopt a centralized assumption for determining optimal ALOHA parameters, wherein the network computes and disseminates the settings rather than vehicles making decentralized decisions. This modeling choice is justified by the stringent reliability and latency requirements of automotive radar systems, where distributed contention-based optimization would introduce instability and risk violating the end-to-end latency budgets mandated by V2X safety applications (e.g., <5<5 ms in 3GPP Rel-18 [3gpp_rel18_v2x]). As vehicles increasingly operate as 5G/6G UEs, network entities such as gNBs or RSUs can act as controllers that compute globally optimal radar access parameters and broadcast them via system information blocks with negligible overhead, thereby supporting ultra-reliable low-latency communications (URLLC). This network-assisted dissemination is consistent with ongoing NR-V2X standardization efforts [3gpp_rel18_v2x], whereas no standardized MAC protocol currently exists for automotive radars, with existing solutions remaining vendor-specific and at the pre-standardization research stage. Since our focus is on the characterization of the tracking probability, we assume a simplistic block ALOHA protocol. Other MAC schemes such as CSMA would require the development of hard-core point processes, e.g., see [alfano2012new], which we will treat in a future work.
Let the time frames be divided into blocks of $T$ slots each. In each block, only a subset of the vehicles are allowed to activate their radars. The activation set follows the classical ALOHA protocol with parameter $\delta\in[0,1]$. Thus, each vehicle in $\Phi_{\rm P}$ operates in a block with probability $\delta$. Within the block in which it is active, it attempts to detect the target in each of the $T$ slots. We present our analysis from the perspective of a single such block in which the ego radar is active. Later, we discuss the intricacies of designing the block length in Section IV. Here we define the key metrics that we statistically characterize in the next section.
@@
We refer to $\nu$ as the reliability parameter of the radar application. The higher the value of $\nu$, more stringent is the requirement from the radar system. The overall protocol is illustrated in Fig. 2. Due to the block ALOHA access policy, the radar gets access in only a fraction of the blocks. In the active blocks, the radar attempts to detect the target in each of the $T$ slots, i.e., it transmits on all the slots of the active block. Each slot corresponds to one radar pulse as shown. The success of a slot is defined in Def. 2, while the target is termed to be successfully tracked if at least $\nu$ consecutive slots are successful. An illustration with $\nu=5$ is shown.
The key challenge in characterizing the above is the fact that the bursts are not independent of each other. To illustrate, consider the event $E_{[t_{1},t_{2}]}$ that all the slots from $t_{1}$ to $t_{2}$ of the block were successful (i.e., the target was successfully detected). Let $\bar{E}_{[t_{1},t_{2}]}$ represent the complement of the event. Naturally, assuming $\nu=4$, $\mathbb{P}\left(E_{[2,5]}|\bar{E}_{[1,4]}\right)=0\neq\mathbb{P}\left(E_{[2,5]}\right)$. Indeed given that the first four slots have failed, the probability that the slots 2 to 5 has a successful run length is 0 due to the slots $2,3,$ and $4$ being common to both the events.
@@ III Statistics of Radar Tracking
The two constituent terms of the conditional detection probability correspond to the impact of the noise and the impact of the interference on the detection performance of the ego radar. Naturally, for a noise-limited regime, the second term of the above should be considered 1, while for an interference-dominated regime, the first term should be considered 1. This is mathematically stated as
$P^{\text{Noise limited}}_{\rm d}$ $={\exp{\left(\frac{-\gamma}{\xi_{0}R^{-2\alpha}}\right)}}$
$P^{\text{Interference dominated}}_{\rm d}$ $=\prod_{{\bf w}\in\phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}$ (5)
Remark 2.
The conditional success probability in (4) is different from that derived in [ghatak2022radar, haritha2021slotted, sukumaran2022slotted]. In particular, the above works consider random access in each slot and hence, there the transmission state of each radar needs to be averaged out with respect to the ALOHA parameter. However, in our case, once the transmission states of the radars are determined at the beginning of the block, the same remains fixed throughout all the slots of the block. Hence, the impact of the ALOHA only appears in the set over which the product of the point functions is evaluated in (4).
In the above, we assumed Rayleigh fading for the interfering signals. The framework can be extended to account for log-normal shadowing by introducing a multiplicative shadowing factor $X_{w}\sim\text{Lognormal}(0,\sigma^{2})$ on each interfering link. In this case, the interference from radar $w$ becomes
[…]
where $h_{w}$ denotes the small-scale fading component. This is elaborated on the following remark.
Remark 3.
In case the interfering links undergo large-scale shadow fading modeled using a log-normal distribution, the conditional detection probability is modified as
$P^{\rm Shadow}_{\rm d}$ $=\mathbb{E}\left[\exp{\left(\frac{-\gamma}{\xi_{0}R^{-2\alpha}}\right)}\left(\prod_{{\bf w}\in\phi^{\prime}}\frac{1}{1+\frac{\gamma X_{\bf w}d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right],$
$=\exp{\left(\frac{-\gamma}{\xi_{0}R^{-2\alpha}}\right)}\left(\prod_{{\bf w}\in\phi^{\prime}}\mathbb{E}\left[\frac{1}{1+\frac{\gamma Xd_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right]\right)$ (6)
where the expectation is with respect to the log-normal random variable $X$. Since no closed-form exists for the log-normal case, this expectation can be evaluated numerically (e.g., via Gauss-Hermite quadrature [liu1994note]). The rest of the analysis, including the run-length distribution remains valid with this modified interference term.
@@
Theorem 1.
Leveraging de Moivre’s solution [hald2005history, Section 22.6], the conditional complementary cumulative density function (CCDF) ${F}_{L}(\nu)=\mathbb{P}\left(L\geq\nu\mid\Phi\right)$ of $L$ as a function of $P_{\rm d}$ is
@@
Given that the ego radar is active in a block, the moments of conditional success probability $P_{\rm d}$ of the ego-radar at a given time $t$ within the block is given by
@@
$=\exp{\left(\frac{-l\gamma}{\xi_{0}R^{-2\alpha}}\right)}\exp\left(-\lambda\delta\int_{0}^{\infty}1-\mathcal{T}_{z}^{l}{\rm d}z\right)\cdot$
$\exp\left(-2\lambda_{L}\int_{0}^{2\pi}\int_{0}^{\infty}1-\exp\left(-\delta\lambda\int_{a_{i}}^{b_{i}}1-\mathcal{T}^{l}_{z,x}{\rm d}z\right){\rm d}x{\rm d}\theta\right),$ (8)
where $\mathcal{T}_{z}=\left(\frac{R^{-2\alpha}}{R^{-2\alpha}+\gamma z^{-\alpha}}\right)$, $\mathcal{T}_{z,x}=\frac{R^{-2\alpha}}{R^{-2\alpha}+\gamma w(x,z,\theta))^{-\alpha}}$ and $w^{2}(x,z,\theta)=x^{2}+z^{2}+2xz\cos\theta$.
@@
It is easy to see that the expression contains a finite number of terms. For example, for a block length of $T=19$ slots and $\nu=9$ slots, $\lfloor\frac{T+1}{\nu+1}\rfloor=2$. Accordingly, the number of total terms in the summation is 8. Recall that the conditional detection probability $P_{\rm d}$ is upper bounded by 1 (since it is a probability). Thus, $P_{\rm d}^{l}\leq 1$ and consequently, due to monotonicity of expectations, we have $\mathbb{E}[P^{l}_{\rm d}]=\zeta(k)\leq 1$. We complete the argument by stating that since the expression of Theorem 9 consists of a finite number of bounded terms, the expression for $P_{\rm t}$ converges trivially.
@@ IV QoS-Aware Optimal Frame Design
In this section, we use the above framework for the design of the transmit frames and the selection of optimal access probability from the perspective of specific automotive use-cases. Let us define $T_{\rm L}$ as the latency requirement specified by the use-case. As an example, low-latency applications such as lane change assistance and automatic emergency braking (AEB) may employ a $T_{\rm L}$ of the order of 1 ms (note that this not the end-to-end latency, but only the latency over the radio link). On the contrary, use-cases that have less stringent latency requirements such as blind spot detection, park assist etc have $T_{\rm L}$ of the order of 10 ms. Some indicative latency requirements are outlined in Table II. Furthermore, we introduce a reliability measure based on the parameter $\nu$. in particular, a larger value of $\nu$ corresponds to a higher reliability requirement from the service. For the tracking service to be rendered successful, let us assume that within the latency deadline, the use-case necessitates at least one successful tracking event, i.e., the ego radar experiences at least one burst of $\nu$ consecutive successes within a deadline of $T_{\rm L}$.
@@ V Numerical Results and Discussion
In this section we discuss the salient features of the network with the help of some numerical results. The radar parameters used for the numerical results are taken from [series2014systems]. Specifically, $P=10$ dBm, $\bar{\sigma}=30$ dBsm, $\alpha=2$, $G_{\rm t}=10$ dBi, $f_{c}=76.5$ GHz, and $N_{0}=-174$ dBm/Hz. Unless otherwise stated, we assume $R=15$ m. The street and vehicle densities are from [shah2024modeling] where the authors have reported real-world data on the same.
@@ V-A Detection vs Tracking Performance
Figure 4: Impact of shadowing on the trends of the tracking probability with respect to $\nu$.
V-B Impact of Shadowing and System Parameters
Fig. 4 shows that log-normal shadow fading deteriorates the interference signals thereby improving the tracking probability. In general, the ego-radar intends to track the vehicle that is directly in front or in the rear of it. Accordingly, it is a reasonable assumption that the target vehicle is in line-of-sight. Thus, shadowing improves the tracking performance without deteriorating the useful reflected signal.
Figure 5: Impact of the system parameters on the tracking probability with respect to $\nu$. Here $R=10$ m.
To illustrate the effect of key parameters on the tracking probability, Fig. 5 plots the probability of maintaining a run of length $\nu$ under different values of the path-loss exponent $\alpha$, beamwidth $\Omega$, and frame length $T$. We observe that increasing the beamwidth (e.g., from $\Omega=30^{\circ}$ to $45^{\circ}$) significantly reduces the tracking probability, as the wider beam admits more interferers. Conversely, increasing the path-loss exponent (from $\alpha=2.0$ to $\alpha=3.5$) improves the tracking probability by attenuating the interference more strongly - the impact on the two-way reflect signal is lower due to the low target distance. Finally, increasing the frame length $T$ enhances the probability of sustaining longer runs, thereby yielding better tracking performance. These trends confirm the sensitivity of radar tracking performance to physical-layer and system design parameters, and highlight the importance of careful parameter tuning.
V-C Impact of Street and Vehicle Density
@@
Figure 8: Expected number of bursts with a length $\geq\nu=8$ for $T=20$ for different $\lambda_{P}$ and $\lambda_{L}$.
@@
TABLE II: Use cases and corresponding latency and reliability requirement derived from [boban2018connected] and [3gpp2019service]. However, note that the latency requirement mentioned here is not end-to-end latency, but the latency over the radio link. The figures are indicative and real values are system dependent. The optimal channel access probability is calculated against each use-case. We assume $\lambda=5e-2$ m-1 and a beamwidth of 10 degrees.
V-D Impact of Use-Case QoS Requirements
@@
[…]
V-E Optimal Frame Design
Figure 10: Impact of the frame design policy on the tracking performance. Here we use real data fitting of New Delhi city to extract the PLCP parameter $\lambda_{L}=0.004$ m-1 (see [shah2024modeling] for details) and use a vehicle density of $\lambda=5e-2$ m-1 and $\Omega=10$ degrees.
@@ V-E Optimal Frame Design
Figure 12: Trends of the detection probability for homogeneous PLCP vs inhomogeneous PLCP and BLCP with respect to the distance of the ego-radar from the center of the city.
Figure 13: Trends of the tracking probability for homogeneous PLCP vs inhomogeneous PLCP and BLCP with respect to the distance of the ego-radar from the center of the city for (a) fixed target distance and (b) stochastic target distance.
V-F Impact of Non-Stationarity
Street systems are denser near the city center while becoming sparse in the outskirts. One such non-stationary aspect of street networks was recently taken up in [shah2024binomial] where a new model called BLCP was introduced which models inhomogeneous street networks across different cities. Remarkably, the model was validated using real data collected from maps in [11075590]. Unlike the PLCP model, in an inhomogeneous process, there does not exist a typical point from the perspective of which the performance can be characterized. On the contrary, the performance is bench-marked for a test point conditioned on its location (say $r_{0}$ from the origin). As suggested by the reviewer, we investigate the change in the trends of the tracking probability as the model becomes increasingly non-stationary. Additionally, considering that the vehicular nodes too become sparser as we move out of the city, we assume that the ego-vehicle intends to detect its nearest target. Consequently, we model the location of the target as the nearest point of the doubly stochastic Cox processes.
In Fig. 12, we show the trends of the detection probability for $r_{0}$ for three new Cox process models, the newly introduced BLCP model (e.g., see [shah2024binomial]) and contrast it with the homogeneous PLCP. Among the contending models, the E1-PLCP model is constructed by thinning the homogeneous PLCP points based on their distance from the origin using an exponential decay function, $f(x)=\exp{(-\delta x)}$, where $x$ represents the point’s distance from the origin and $\delta$ is the decay factor. Similarly, in the E2-PLCP model, the intensity of the line process is varied as per an exponential-based intensity function, $\lambda(r)=\exp{(-\delta r)}$, where $r$ is the perpendicular distance of a PLCP line from the origin. In the E-BLCP model, an additional layer of randomness is added to the BLCP process by generating non-homogeneous PPPs on each line with intensity function $\lambda(x)=\exp{(-\delta x)}$, where $x$ denotes the distance of a BLCP point from the generating point of its respective line. This results in a more realistic distribution of points concentrated around the city center, enhancing the realism of the original BLCP model. Fig. 12 shows that for BLCP and E-BLCP, the detection success probability initially rises with $r_{0}$, peaks, and then decreases. The point of minimum value of $\mathbb{E}\left[\frac{R}{d_{1}}\right]$, where $R$ is the target distance and $d_{1}$ is the closest interferer distance is the same as the value where success probability is maximum. It is worth recalling here that now both $d_{1}$ and $R$ are random quantities. In the case of homogeneous PLCP, naturally, the probability of success is constant at all locations.
We extend this discussion to the tracking probability in Fig. 13, where we plot the cases of fixed target distance (Fig. 13a) and random target distance (Fig. 13b). Here for the BLCP model, we have assumed the radius of the generating circle as 100 m. For the fixed target distance, the we see that the tracking probability increases as the ego-radar goes out of the city center and towards the outskirts. This is due to the decrease in interference as a result of less dense presence of the vehicles as $r_{0}$ increases. However, for the random target distance, as the vehicles become less dense, not only the interference decreases but the target also moves statistically farther from the ego-radar for the E1-PLCP and E-BLCP models, and hence we note a deterioration of the tracking probability. In a realistic deployment, the ego-radar can be equipped with a change-detection module (e.g., see [ghatak2021kolmogorov]) to sense the vehicle density and adapt its radar parameters accordingly.
V-G Implementation Considerations
TABLE III: Execution time benchmarks for evaluating the analytical expressions.
Implementation Average runtime per evaluation Relative to 20 ms frame
MATLAB (prototype) 0.40 ms 2% of frame
Python (NumPy/SciPy) 0.30 ms 1.5% of frame
Optimized C (single-core) 0.18 ms 0.9% of frame
Optimized C (multi-core, 4 threads) 0.08 ms 0.4% of frame
Real-time feasibility benchmarks: It is worth to highlight that the implementation of the protocol proposed in this work involves solving multiple integrals and summation with on-board processors in automotive systems, which can be challenging. However, with recent advancements of radar processing units appended with hardware-accelerated processing (e.g., see the discussion in  [engels2021automotive, meinl2017real]), the implementation of such protocols are feasible. To substantiate our claim, Table III reports the average execution time of a single evaluation of the Laplace transform of the interference distribution and the required summations, obtained using MATLAB, Python, and optimized C implementations. All tests were carried out on an Intel i7-1165G7, 2.8 GHz CPU with 16 GB RAM. The results indicate that our framework is executable within a fraction of the radar frame duration (10–50 ms), leaving sufficient computational margin. These results confirm that the proposed computations are well within real-time feasibility, even without hardware acceleration. Further improvements are possible using pre-computed lookup tables or FPGA/GPU offloading.
On centralized MAC: Additionally, the proposed scheme consists of the derivation of an optimal ALOHA parameter which is to be calculated in the network operator for the connected cars and broadcast through network to vehicle (N2V) links. Furthermore, the vehicle density across different times of the day and across different locations of a city albeit varying, remain relatively stable in the order of an hour. In this regard, similar to how a network operator sets the cell-range expansion parameters in a network of base stations, it is envisaged that the network operator will approximate the current vehicular density and fix the optimal ALOHA parameter for several hours. Indeed this is how most of macro network parameters are selected in current deployment of networks. From a systems implementation perspective, such network-assisted parameter broadcasting eliminates the instability associated with distributed contention-based optimization. Importantly, it allows the network to dynamically adapt ALOHA parameters to spatio-temporal traffic variations and heterogeneous service requirements across vehicles. The design is thus aligned with the ongoing 3GPP NR-V2X standardization efforts, where the cellular network plays a key role in providing control-plane assistance for both communication and sensing functions [3gpp_rel18_v2x].
Comparison to other MAC protocols: Existing MAC protocols, such as time-division multiplexing or contention-based schemes, primarily focus on coordinating medium access in a heuristic manner. While these methods are straightforward to implement, they lack analytical characterizations of performance under stochastic interference. In contrast, our proposed framework provides closed-form analytical guarantees on the tracking probability, offering insights into the fundamental limits of radar coexistence. A direct quantitative comparison is challenging because MAC-level heuristics and our analytical model address complementary aspects of the problem. Nevertheless, our results suggest that by integrating protocol-level scheduling with the interference-aware run-length analysis developed here, one could design MAC protocols that are both analytically tractable and practically efficient. We identify this integration as an important direction for future work.
@@ Appendix A Proof of Lemma 1
We derive the expressions of $a$ and $b$ using simple geometrical arguments from Fig. 14, which corresponds to the scenario presented in Fig. 1. the ego-radar is shown in green at the origin $O$, while the interfering radar is located at $A$ and is shown in red. The interfering radar is present in an intersecting street $L_{0}$ which is is affine extension of the segment $BE$. The two key angles of interest are $\angle AOG=\psi_{1}$ and $\angle OBG=\psi_{2}$. Recall that the half-power beamwidth of the radar is $\Omega$. Hence the angle $\angle OAE=\frac{\Omega}{2}$. The interfering radar in this illustration is shown at a distance $AG=a$ from the intersection point. Now, employing the sine rule in the triangle $AOG$, we have:
$\frac{a}{\sin\Psi_{1}}=\frac{d}{\sin\frac{\Omega}{2}}.$
Now since the orientation of the intersecting line is $\theta$ with respect to the $x$-axis, $\angle OEA=\frac{\pi}{2}-\theta$. This implies $\angle OGE=\theta$ and accordingly, $\angle OGA=\pi-\theta$. Consequently, we see from the triangle OAG that $\psi_{1}=\pi-\angle OAG-\angle OGA=\pi-\frac{\Omega}{2}-(\pi-\theta)=\theta-\frac{\Omega}{2}$. Combining this with the sine rule, we have
$a=\frac{d\sin\psi_{1}}{\sin\frac{\Omega}{2}}=\frac{d\sin\left(\theta-\frac{\Omega}{2}\right)}{\sin\frac{\Omega}{2}}.$ (11)
The value of $\Psi_{2}$ and accordingly $b$ follows similarly.